# CroMe: Multimodal Fake News Detection using Cross-Modal Tri-Transformer and Metric Learning

### v1

@@ Abstract
Multimodal Fake News Detection has received increasing attention recently. Existing methods rely on independently encoded unimodal data and overlook the advantages of capturing intra-modality relationships and integrating inter-modal similarities using advanced techniques. To address these issues, Cross-Modal Tri-Transformer and Metric Learning for Multimodal Fake News Detection (CroMe) is proposed. CroMe utilizes Bootstrapping Language-Image Pre-training with Frozen Image Encoders and Large Language Models (BLIP2) as encoders to capture detailed text, image and combined image-text representations. The metric learning module employs a proxy anchor method to capture intra-modality relationships while the feature fusion module uses a Cross-Modal and Tri-Transformer for effective integration. The final fake news detector processes the fused features through a classifier to predict the authenticity of the content. Experiments on datasets show that CroMe excels in multimodal fake news detection.
1 Introduction
@@
Advances in information technology and the growing use of social media have made these platforms central to information exchange Mitra et al. (2017). However, this has also led to a rise in fake news, which harms public opinion, disrupts political stability Shu et al. (2017), and affects social and economic activities Xia et al. (2023). Detecting false information is essential to prevent its spread. The shift to digital platforms has changed how people consume news, increasing exposure to misinformation. Traditional methods, such as identifying logical flaws or obvious signs like spelling errors and image alterations, were effective for text or image content. However, the rise of multimedia formats like images and videos has accelerated the spread of fake news Jin et al. (2017), making multimedia-focused detection methods necessary.
Figure 1 shows four examples from the Weibo and Weibo-21 datasets. (a) A group of Quora users listed historically ingenious but unreasonable weapons which are semantically contradictory. (b) In Marco Island, dolphins rescued a dog from a canal and firefighters reunited it with its owner which are semantically conflicting. (c) Yichang police seized 7 macaques and arrested 2 suspects which is unrelated to the tiger in the image. (d) The text mentions abandoned Douglas DC 3 of the Aurora’s light but the image shows only the Aurora’s light. The main challenge in multimodal fake news detection is effectively identifying incongruent semantic features in both intra-modality and inter-modality contexts.
@@
Figure 1: Some fake examples from the Weibo and Weibo-21 datasets include: (a) The image shows intra-modality inconsistency. (b) Both the image and text contain intra-modality inconsistencies. (c) The image and text are unrelated. (d) The image does not match the text.
Fake news detection has progressed from text-based methods to advanced deep learning techniques. Early work focused on text analysis Castillo et al. (2011), while later studies utilized deep neural networks for linguistic and temporal patterns Ma et al. (2016), and attention mechanisms with RNNs Chen et al. (2018). Jin et al. Jin et al. (2016) introduced multimodal approaches by combining image, text, and social context features. Choi et al. Choi and Kim (2024) developed TT-BLIP, integrating text, image, and multimodal features using BLIP Li et al. (2022) encoders and a MultiModal Tri-Transformer for feature fusion. However, it primarily focuses on modality combination without addressing inter-modal and intra-modal relationships.
This paper introduces a Multimodal Fake News Detection approach using the Cross-Modal Tri-Transformer and Metric Learning (CroMe) network, a motivated framework inspired by the TT-BLIP model. Text is processed using pre-trained BERT Devlin et al. (2019), and images are analyzed with Masked Autoencoders (MAE) He et al. (2022). The BLIP2 Li et al. (2023) model captures semantic features across modalities. Additionally, CroMe introduces Cross-Modal fusion Chen et al. (2022) to compute and integrate cross-modal similarities, addressing the limitations of existing methods in capturing complex inter-modal interactions. For intra-modal relationships, a type of metric learning Schroff et al. (2015); Movshovitz-Attias et al. (2017) called the Proxy Anchor Loss method Kim et al. (2020) is used. This method captures fine-grained similarities while maintaining the efficiency of proxy-based methods through dynamic gradients. Thus, CroMe ensures that features within the same modality are more closely aligned while capturing and integrating cross-modal correlations, enhancing the overall feature representation for more accurate fake news detection. Experiments performed using the Weibo and Weibo-21 datasets demonstrate that CroMe outperforms state-of-the-art models, highlighting its effectiveness in Multimodal Fake News Detection. For the Politifact dataset, CroMe achieved results comparable to state-of-the-art models due to its smaller size.
This paper offers four key contributions:
  • The proposed multimodal fake news detection model, called CroMe, utilizes the advanced BLIP2 model for feature extraction to capture detailed text, image, and combined image-text features.
@@
  • CroMe is evaluated using the Weibo and Weibo-21 datasets, demonstrating that it outperforms state-of-the-art models and highlights its effectiveness in multimodal fake news detection. For the Politifact dataset, CroMe achieved comparable results to state-of-the-art models, likely due to the dataset’s smaller size.
2 Related work
Figure 2: Overview of the CroMe architecture. Masked Autoencoders (MAE), BERT, and BLIP2 encode multimodal features. Metric Learning extracts intra-modal relationships by representing class data points with proxies, where arrow thickness indicates the gradient scale. The CMTTF and Fake News Detector modules use cross-modal fusion and fake news detection.
Several methods for multimodal fake news detection focus on extracting features from text and images. EANN Wang et al. (2018) uses an event discriminator, and MCAN Wu et al. (2021) integrates textual and visual features through co-attention layers. MVAE Khattar et al. (2019) employs a multimodal variational autoencoder to reconstruct text and image features. Spotfake Singhal et al. (2019) applies BERT Devlin et al. (2019) for text and VGG19 Simonyan and Zisserman (2014) for images, with Spotfake+ Singhal et al. (2020) extending this to full articles. SAFE Zhou et al. (2020) detects fake news by analyzing text and visual similarities, while CAFE Chen et al. (2022) measures cross-modal ambiguity using KL divergence Kullback and Leibler (1951). LIIMR Singhal et al. (2022) emphasizes primary modalities and minimizes less significant ones. DistilBert Allein et al. (2021) detects disinformation by analyzing user preferences and sharing behaviors. BDANN Zhang et al. (2020) combines multimodal features to reduce event-specific biases. FND-CLIP Zhou et al. (2023) uses the CLIP model to measure image-text correlations and aggregate features through modality-wise attention. TT-BLIP Choi and Kim (2024) employs BLIP Li et al. (2022) encoders for text, image, and multimodal data, with the Multimodal Tri-Transformer fusing features using multi-head attention mechanisms.
Unlike previous models, the proposed model integrates uni-modal features and cross-modal correlations. The metric learning module aligns features within the same modality, while cross-modal similarities capture inter-modality interactions. By focusing on both intra-modality alignment and inter-modality integration, this approach improves fake news detection in social media.
3 Methodology
@@
3.1 Overview
CroMe, introduced in this paper, as illustrated in Figure 2, comprises four modules: the encoder, metric learning, Cross-Modal Tri-Transformer Fusion (CMTTF), and fake news detection. The encoder extracts image, text, and integrated encodings. Metric learning enhances intra-modality relationships using the Proxy Anchor method. The CMTTF module fuses these encodings, and the detection module determines whether the content is real or fake.
@@ 3.1 Overview
3.2 Encoder Module
The encoder layer consists of components for encoding text, images and their combination. These components process input data denoted as $x_{\text{img}}$ for images and $x_{\text{txt}}$ for text.
The image encoder utilizes two parallel methods: Masked Autoencoders (MAE) He et al. (2022) and BLIP2 Image. The pretrained MAE functions as the primary image encoder, producing encodings $Z_{\text{i1}}=f_{\text{MAE}}(x_{\text{img}})$, capturing both global and local features by reconstructing missing data parts. In parallel, BLIP2 Image provides an alternative encoding $Z_{\text{i2}}=f_{\text{BLIP2\_i}}(x_{\text{img}},x_{\text{txt}})$, ensuring that image-specific encodings are derived without textual influence by utilizing a “dummy text” input.
For textual data, BERT and BLIP2 Text encoders are utilized. The pretrained BERT Devlin et al. (2019) generates encodings $Z_{\text{t1}}=f_{\text{BERT}}(x_{\text{txt}})$ by using its bidirectional understanding of text context. Additionally, BLIP2 Text Li et al. (2023) produces encodings $Z_{\text{t2}}=f_{\text{BLIP2\_t}}(x_{\text{img}},x_{\text{txt}})$ by focusing on text data alone, using a zero-filled image tensor to isolate the textual encoding process. The outputs from these encoding processes are concatenated to form integrated image and text encodings, denoted as $Z_{\text{i}}$ and $Z_{\text{t}}$, respectively, for use in subsequent model components.
The BLIP2 model generates combined Image-Text encodings, capturing the cross-modal information between these modalities. This is represented as $Z_{\text{b}}=f_{\text{BLIP2}}(x_{\text{img}},x_{\text{txt}})$, incorporating the integrated features from both image and text inputs.
3.3 Metric Learning
@@
Proxy anchor loss Kim et al. (2020) effectively learns data representations by utilizing distance relationships between data points and proxies. In this work, it captures intra-modal relationships across different modalities, improving the distinction between rumors and non-rumors. The loss function $L(X)$ uses static proxy assignment Movshovitz-Attias et al. (2017), selecting a data point from each class as a proxy. It minimizes the distance between the proxy and same-class data points while maximizing the distance to those from different classes. Modality embeddings $X$ are split into positive ($X^{+}_{p}$) and negative ($X^{-}_{p}$) sets relative to their proxies. The loss is defined as:
$\begin{split}L(X)=&\frac{1}{|P^{+}|}\sum_{p\in P^{+}}\log\bigg{(}1+\sum_{x\in X^{+}_{p}}e^{-\alpha(s(x,p)-\delta)}\bigg{)}\\ &+\frac{1}{|P|}\sum_{p\in P}\log\bigg{(}1+\sum_{x\in X^{-}_{p}}e^{\alpha(s(x,p)+\delta)}\bigg{)}\,,\end{split}$
@@
As shown in Figure 2, the model iteratively trains each modality $*Z_{\text{*}}$ by fixing the parameters of other modalities, allowing it to refine intra-modal relationships without interference. This process is applied to each $Z_{\text{*}}\in\{Z_{\text{i1}},Z_{\text{i2}},Z_{\text{t1}},Z_{\text{t2}},Z_{\text{b}}\}$. Proxy Anchor Loss, illustrated in Figure 3(a), adjusts pull and push forces based on data point proximity, unlike traditional triplet loss Schroff et al. (2015); Peng et al. (2023). Proxy Anchor Loss is computationally efficient as it calculates distances between proxies and data points rather than between individual points, reducing computational load. Applied to text, image, and image-text features, it clusters data points of the same class while separating those of different classes, enhancing fake news detection.
3.4 Cross-Modal Tri-Transformer Fusion
This section introduces Cross-Modal Tri-Transformer Fusion (CMTTF) for integrating and processing image, text, and image-text data. CMTTF combines Cross-Modal Fusion Chen et al. (2022) and Tri-Transformer Choi and Kim (2024), as illustrated in Figure 4(a), with details on each component provided below.
Cross-Modal Fusion
To capture semantic interactions between modalities, the Cross-Modal Fusion module integrates text-image similarities using dot product with softmax normalization and cosine similarity Luo et al. (2018). Given unimodal representations $Z_{\text{i1}}$, $Z_{\text{t1}}$, and additional representations $Z_{\text{i2}}$, $Z_{\text{t2}}$, $Z_{\text{b}}$, the process follows these steps: 1) Inter-modal Fusion and 2) Cosine Similarity.
Inter-modal Fusion integrates semantic interactions between text and image modalities (is illustrated in Figure 4(b)). Correlations between text features $\mathbf{t}$ and image features $\mathbf{i}$ are computed, normalized, and combined into a unified representation $\mathbf{C}_{\text{1}}$, as shown:
$\mathbf{f}_{t\rightarrow i}=\text{Softmax}\left(\frac{\mathbf{t}\cdot\mathbf{i}^{\mathbf{T}}}{\sqrt{d}}\right),$ $\;\mathbf{f}_{i\rightarrow t}=\text{Softmax}\left(\frac{\mathbf{i}\cdot\mathbf{t}^{\mathbf{T}}}{\sqrt{d}}\right),$
$\mathbf{C}_{\text{1}}=\mathbf{f}_{t\rightarrow i}$ $\oplus\mathbf{f}_{i\rightarrow t}\,.$

### v2

# This work was supported by the government of the Republic of Korea (MSIT) and the National Research Foundation of Korea (RS-2024-00466496), and supported by Korea Industrial Technology Association (KOITA) grant funded by Korea Ministry of Science and ICT (MSIT) (KOITA20250002-62)

@@ Abstract
Multimodal fake news detection has received increasing attention recently. Existing methods rely on independently encoded unimodal data and overlook the advantages of capturing intra-modal relationships and integrating inter-modal similarities using advanced techniques. To address these issues, Cross-Modal Tri-Transformer and Metric Learning (CroMe) for multimodal fake news detection is proposed. CroMe utilizes bootstrapping language-image pre-training (BLIP) with frozen image encoders and large language models as encoders to capture detailed text, image, and combined image-text representations. The metric learning module employs a proxy anchor method to capture intra-modality relationships while the feature fusion module uses a Cross-Modal and Tri-Transformer for effective integration. The final fake news detector processes the fused features through a classifier to predict the authenticity of the content. Experiments on datasets show that CroMe excels in multimodal fake news detection.
Index Terms:
Multimodal fusion, Multimodal learning, Metric learning, Fake news detection
[…]
=-15pt
I Introduction
@@
With rapid advances in information technology and the growing use of social media, digital online platforms have become the center of information exchange [mitra2017parsimonious]. However, this has also led to a significant rise in fake news and misinformation, which may harm public opinion, disrupt political stability, and affect social and economic activities [zhu2022locating, shu2017fake, xia2023covid]. Thus, effective false information detection has become essential to prevent these issues. Traditional methods, such as identifying logical flaws or obvious signs like spelling errors and image alterations, were effective for text or image content separately. However, the rise of multimedia formats like images and videos has accelerated the spread of fake news [jin2017multimodal], making multimedia-focused detection methods necessary.
Figure 1 shows four different examples of fake news from the Weibo [jin2017multimodal] and Weibo-21 datasets [nan2021mdfend], which reveal various types of semantic inconsistencies. The first example is fake news using a manipulated image, which shows intra-modal inconsistency. Next, the second example shows that both the text and image are independently manipulated, demonstrating intra-modality inconsistencies within each modality. In the third example, the tiger image is unrelated to the text, demonstrating inter-modality inconsistency. The last example shows inter-modality conflict, where the text description does not fully align with the image. Therefore, the main challenge in multimodal fake news detection is to effectively identify incongruent semantic features in both intra-modality and inter-modality contexts. Recent advances in fake news detection have evolved from text-based methods to advanced deep learning approaches. Early work focused on text analysis [castillo2011information], while later studies employed deep neural networks (DNNs) for linguistic and temporal patterns [ma2016detecting], and applied attention mechanisms with recurrent neural networks (RNNs) [chen2018call]. Jin et al. [jin2016novel] introduced multimodal approaches by combining image, text, and social context features. More recently, Choi et al. [10706486] developed TT-BLIP, integrating text, image, and multimodal features using BLIP [li2022blip] encoders and a MultiModal Tri-Transformer for feature fusion. Despite these advances, however, these approaches primarily focus on modality combination without addressing inter-modal and intra-modal relationships.
@@
Figure 1: Some fake examples from the Weibo and Weibo-21 datasets include: (a) The image shows intra-modality inconsistency. (b) Both the image and text contain intra-modality inconsistencies. (c) The image and text are unrelated, indicating inter-modality inconsistency. (d) The image does not match the text, reflecting inter-modality conflict.
This paper introduces a multimodal fake news detection approach using the Cross-Modal Tri-Transformer and Metric Learning (CroMe) network, a motivated framework inspired by the TT-BLIP model. To capture detailed text, image, and combined image-text features, CroMe uses pre-trained BERT [devlin-etal-2019-bert] for text encoding, Masked Autoencoders (MAE) [he2022masked] for image feature extraction, and BLIP2-OPT [li2023blip] for semantic features across modalities. CroMe also introduces a Cross-Modal Fusion [chen2022cross] to compute and integrate cross-modal similarities, addressing the limitations of existing methods in capturing complex inter-modal interactions. For intra-modal relationships, a type of metric learning called the Proxy Anchor Loss method [schroff2015facenet, movshovitz2017no, kim2020proxy] is used. This method captures fine-grained similarities while maintaining the efficiency of proxy-based methods through dynamic gradients. Thus, CroMe ensures that features within the same modality are more closely aligned while capturing and integrating cross-modal correlations, enhancing the overall feature representation for more accurate fake news detection. Experiments performed using the Weibo and Weibo-21 datasets demonstrate that CroMe outperforms state-of-the-art models, highlighting its effectiveness in Multimodal Fake News Detection. For the Politifact dataset, CroMe achieved results comparable to state-of-the-art models because of the smaller dataset size. The main contributions of this work are summarized as follows:
  • The proposed multimodal fake news detection model, called CroMe, utilizes the advanced BLIP2-OPT model for feature extraction to capture detailed text, image, and combined image-text features.
@@
The paper is organized as follows. Related work is described in Section II. Section III introduces the proposed CroMe model. Then, Section IV assesses its performance and Section V concludes the paper.
Figure 2: Overview of the CroMe architecture. Masked Autoencoder (MAE), BERT, and BLIP2-OPT encode multimodal features. Metric learning extracts intra-modal relationships by representing class data points with proxies, where arrow thickness indicates the gradient scale. CMTTF and fake news detector modules use cross-modal fusion and fake news detection.
II Related work
Several approaches have been developed for multimodal fake news detection, emphasizing the extraction of features from both images and text. EANN [wang2018eann] enhances feature extraction using an event discriminator. MCAN [wu2021multimodal] integrates textual and visual features through multiple co-attention layers. MVAE [khattar2019mvae] employs a multimodal variational autoencoder to learn latent variables and reconstruct text and image features. Spotfake [singhal2019spotfake] applies BERT [devlin-etal-2019-bert] for text and VGG19 [simonyan2014very] for image features, with Spotfake+ [singhal2020spotfake+] extending this to full articles. SAFE [zhou2020similarity] uses a similarity-based method to detect fake news by analyzing text and visual information, and CAFE [chen2022cross] measures cross-modal ambiguity by calculating the Kullback-Leibler (KL) divergence [kullback1951information] between unimodal feature distributions. LIIMR [singhal2022leveraging] emphasizes the primary modality while minimizing the influence of less significant ones. DistilBert [allein2021like] detects disinformation by correlating user preferences and sharing behaviors using latent representations of content. BDANN [zhang2020bdann] combines features from two modalities to tackle event-specific biases in fake news detection on microblogs. FND-CLIP [zhou2023multimodal] uses the CLIP vision-language model to measure image-text correlations, employing unimodal and pairwise CLIP encoders to aggregate features through modality-wise attention. TT-BLIP[10706486] improves direction by using the BLIP[li2022blip] model for unified vision-language understanding, with separate encoders for text, images, and multimodal data. The Multimodal Tri-Transformer then fuses these features using multi-head attention mechanisms for better data analysis.
Unlike previous models, the proposed model integrates unimodal features and cross-modal correlations. The metric learning module focuses on intra-modality relationships, ensuring that features from the same modality are closely aligned. This paper also captures inter-modality interactions by computing and integrating cross-modal similarities. By focusing on both intra-modality alignment and inter-modality integration, this approach improves fake news detection and outperforms the performance of state-of-the-art multimodal detection models using benchmark datasets.
III Methodology
@@
III-A Overview
CroMe, introduced in this paper and illustrated in Figure 2, is a multimodal model for fake news detection comprising four key modules. First, the encoder module extracts image encodings, text encodings, and integrated image-text encodings to capture modality-specific features from the input data. Second, the metric learning module enhances intra-modality relationships using the proxy anchor method [kim2020proxy], which effectively clusters same-class data points while separating different-class points. Third, the Cross-Modal Tri-Transformer Fusion (CMTTF) module fuses the modality-specific features generated by the encoder to capture complex cross-modal interactions and semantic relationships. Both the metric learning and CMTTF modules are applied in parallel to the encoder outputs, and their respective losses are combined to form the total loss $L_{\text{total}}$. This total loss is used to update the encoder parameters in the subsequent training epoch, enabling the encoder to capture both intra-modal and inter-modal relationships. Lastly, the fake news detection module determines whether the news is real or fake from the fused representation. This four-module architecture enables CroMe to effectively capture both modal-specific features and complex cross-modal interactions, resulting in more accurate fake news detection.
III-B Encoder Modules
The encoder layer consists of components for encoding text, images and their combination. These components process input data denoted as $x_{\text{img}}$ for images and $x_{\text{txt}}$ for text. For image data, the image encoder utilizes two parallel methods: Masked Autoencoders (MAE) [he2022masked] and BLIP2-OPT Image. The pretrained MAE functions as the primary image encoder, producing encodings $Z_{\text{i1}}=f_{\text{MAE}}(x_{\text{img}})$, capturing both global and local features by reconstructing missing data parts. In parallel, BLIP2-OPT Image provides an alternative encoding $Z_{\text{i2}}=f_{\text{BLIP2-OPT}_{\text{img}}}(x_{\text{img}},x_{\text{txt}})$, ensuring that image-specific encodings are derived without textual influence by utilizing a “dummy text” input. MAE focuses on global image structures, while BLIP2-OPT Image captures fine-grained semantics and multimodal associations, complementing each other. For textual data, BERT and BLIP2-OPT Text encoders are utilized. The pretrained BERT [devlin-etal-2019-bert] generates encodings $Z_{\text{t1}}=f_{\text{BERT}}(x_{\text{txt}})$ by using its bidirectional understanding of text context. Additionally, BLIP2-OPT Text [li2023blip] produces encodings $Z_{\text{t2}}=f_{\text{BLIP2-OPT}_{\text{txt}}}(x_{\text{img}},x_{\text{txt}})$ by focusing on text data alone, using a zero-filled image tensor to isolate the textual encoding process. While BERT excels in understanding text context, it lacks multimodal capabilities. In contrast, BLIP2-OPT Text, trained on multimodal datasets, is better suited for contexts. The outputs from these encoding processes are concatenated to form integrated image and text encodings, denoted as $Z_{\text{i}}$ and $Z_{\text{t}}$, respectively, for use in subsequent model components. The BLIP2-OPT model generates combined Image-Text encodings, capturing the cross-modal information between these modalities. This is represented as $Z_{\text{b}}=f_{\text{BLIP2-OPT}_{\text{img,txt}}}(x_{\text{img}},x_{\text{txt}})$, incorporating the integrated features from both image and text inputs.
III-C Metric Learning
@@
Proxy anchor loss [kim2020proxy] effectively learns data representations by utilizing distance relationships between data points and proxies. In this work, it is used to capture intra-modal relationships for different modalities, improving the distinction between data points in same class. The loss function $L(X)$ uses static proxy assignment [movshovitz2017no], selecting a data point from each class as a proxy. It minimizes the distance between the proxy and same-class data points while maximizing the distance to those from different classes. Modality embeddings $X$ are split into positive ($X^{+}_{p}$) and negative ($X^{-}_{p}$) sets relative to their proxies. The loss is defined as:
[…]
@@
As shown in Figure 2, the model iteratively trains each modality $Z_{\text{*}}$ by fixing the parameters of other modalities, allowing it to refine intra-modal relationships without interference. This process is applied to each $Z_{\text{*}}\in\{Z_{\text{i1}},Z_{\text{i2}},Z_{\text{t1}},Z_{\text{t2}},Z_{\text{b}}\}$. Proxy Anchor Loss, illustrated in Figure 3(a), adjusts pull and push forces based on data point proximity, unlike traditional triplet loss [schroff2015facenet, peng2023mrml]. Proxy Anchor Loss is computationally efficient as it calculates distances between proxies and data points rather than between individual points, reducing computational load. Applied to text, image, and image-text features, it clusters data points of the same class while separating those of different classes, enhancing fake news detection.
III-D Cross-Modal Tri-Transformer Fusion
This section introduces Cross-Modal Tri-Transformer Fusion (CMTTF) for integrating and processing image, text, and image-text data. CMTTF combines Cross-Modal Fusion [chen2022cross] and Tri-Transformer [10706486], as illustrated in Figure 4(a), with details on each component provided below.
III-D1 Cross-Modal Fusion
To capture semantic interactions between modalities, the Cross-Modal Fusion module integrates text-image similarities using dot product with softmax normalization and cosine similarity [luo2018cosine]. Given unimodal representations $Z_{\text{i1}}$, $Z_{\text{t1}}$, and additional representations $Z_{\text{i2}}$, $Z_{\text{t2}}$, $Z_{\text{b}}$, the process follows these steps: 1) Inter-modal Fusion and 2) Cosine Similarity.
First, Inter-modal Fusion integrates semantic interactions between text and image modalities (is illustrated in Figure 4(b)). Correlations between text features $\mathbf{t}$ and image features $\mathbf{i}$ are computed, normalized, and combined into a unified representation $\mathbf{C}_{\text{1}}$, as shown:
$\mathbf{f}_{t\rightarrow i}$ $=\text{Softmax}\left(\frac{\mathbf{t}\cdot\mathbf{i}^{\mathbf{T}}}{\sqrt{d}}\right),$ (2)
$\mathbf{f}_{i\rightarrow t}$ $=\text{Softmax}\left(\frac{\mathbf{i}\cdot\mathbf{t}^{\mathbf{T}}}{\sqrt{d}}\right),$ (3)
$\mathbf{C}_{\text{1}}$ $=\mathbf{f}_{t\rightarrow i}\oplus\mathbf{f}_{i\rightarrow t}\,.$ (4)